OpenStudy (anonymous):

What is the area of the largest rectangle with lower base on the x-axis and upper vertices on the curve y=12-x^2?

OpenStudy (anonymous):

Let \[ A(x) = 2x(12-x^2)= 24 x - 2x^3\\ A'(x) = 24 - 6 x^2=6( 4-x^2)=6(2-x)(2+x) \] You can verify that A(x) is maximum at x=2 and A(2)=32. The answer is 32